OpenStudy (zolock):

Show all your work. Simplify : x^2 +3x+2 / x+2

OpenStudy (anonymous):

Find the LCD for step #1

OpenStudy (zolock):

is there another way ti do this.

OpenStudy (anonymous):

the first term has a denominator of 1 and the second term has one of 1(x+2)

OpenStudy (anonymous):

since they both have 1 you don't have to do anything with it, but since only the second term has a denominator of (x+2), you have to multiply top and bottom of the first term by (x+2)

OpenStudy (mathstudent55):

First, factor the numerator.

OpenStudy (anonymous):

we are unsure of what the numerator for any term is

OpenStudy (anonymous):

but yes that would be wise

OpenStudy (mathstudent55):

\(\dfrac{x^2 + 3x + 2}{x + 2} \) First, factor the numerator. Try one factor of x + 2, and find the other factor. If you can factor that way, then you can cancel out the common factor. If you can't factor the numerator and get one factor of x + 2, then you can't simplify the fraction because the only factor available in the denominator for reducing is x + 2.

OpenStudy (zolock):

then

OpenStudy (mathstudent55):

Have you factored the numerator?

OpenStudy (zolock):

(x+2)(x+1)

OpenStudy (mathstudent55):

Great. Now let's write that in the fraction. \(\dfrac{x^2 + 3x + 2}{x + 2} \) \(=\dfrac{(x + 2)(x + 1)}{x + 2}\) We are up to here now. Do you see a common factor in the numerator and denominator?

OpenStudy (zolock):

yes

OpenStudy (mathstudent55):

Good. Now you divide both the numerator and the denominator by the common factor.

OpenStudy (mathstudent55):

What do you get?

OpenStudy (zolock):

i get x +1

OpenStudy (mathstudent55):

Correct. \(=\dfrac{\cancel{\color{red}{(x + 2)}}(x + 1)}{\cancel{\color{red}{x + 2}}}\) \(= x + 1 \)

OpenStudy (zolock):

thank you very much mathstudent55

OpenStudy (mathstudent55):

You're welcome.